OpenStudy (anonymous):

in a small bag of candy,there are 9 orange candies and 4 blue candies.you select 1 candy ,eat it,and then select another find P(you select an orange candy and then a blue candy) A. 3/13 B. 36/169 C. 1/36 D. 9/26

OpenStudy (anonymous):

P(O1) = 9/13, P(B2) = 4/12 because there is one less candy in total. This is a sequence so multiply. P(O1 and B2) = 9/13*4/12 = 9/13 * 1/3 = 9/39 = 3/13

OpenStudy (anonymous):

so the answer is 3/13?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thank u very much pilot:)